# non-asymptotic estimates for markov transition matrices with rigorous error bounds

### v1

@@ 3.1.3. Invariant distributions
$\sum\limits_{\bm{v}\in\Omega^{2}}\mu_{2}(\bm{v})p_{2}(\bm{v},\bm{u})$ $=\sum\limits_{v\in\Omega}\mu_{2}\left(v,u_{1}\right)p\left(u_{1},u_{2}\right)$ (3.11)
@@ 3.3. Spectral properties of $\bm{P}_{2}$ and $\tilde{\bm{P}}_{2}$
In order to derive non-asymptotic error bounds for MLE and SCE, we need to study the spectral gaps of $\bm{P}_{2}$ and $\tilde{\bm{P}}_{2}$, respectively. To start, the next theorem focuses on the spectral properties of $\bm{P}_{2}$. We will show that $\bm{P}$ and $\bm{P}_{2}$ share the same nonzero eigenvalues with algebraic multiplicity accounted by decomposing $\bm{P}_{2}$ as the product of two low-rank matrices according to the construction of $\bm{P}_{2}$. Furthermore, in order to obtain the spectral gaps of $\bm{P}_{2}$ required for concentration inequalities, we also study the spectrum of $\bm{P}_{2}\bm{P}_{2}^{*}$ and that of $\bm{P}_{2}+\bm{P}_{2}^{*}$. We find that, the absolute spectral gap of $\bm{P}_{2}$ is always zero, but the symmetric spectral gap of $\bm{P}_{2}$ is at least half of the symmetric spectral gap of $\bm{P}$, which makes it possible to apply certain concentration inequalities to $\bm{P}_{2}$ and derive non-asymptotic tail bounds for MLE.
@@
    $1$ is the only nonzero singular value of the matrix $\bm{D}_{\mu_{2}}^{1/2}\bm{P}_{2}\bm{D}_{\mu_{2}}^{-1/2}$ and the multiplicity of $1$ is $d$. As a result, $\eta_{a}(\bm{P}_{2})\equiv 0$.
@@
    It holds that
    $\eta_{s}(\bm{P}_{2})\geq\frac{\eta_{s}(\bm{P})}{2}>0,$ (3.23)
    and hence, $\mu_{2}$ is the unique invariant distribution of $\bm{P}_{2}$.
@@
    When $\bm{P}$ is reversible, any nonzero eigenvalue of $(\bm{P}_{2}+\bm{P}^{*}_{2})/2$ must be a nonzero eigenvalue of $(\bm{P}+\bm{\mathbf{I}})/2$ or $(\bm{P}-\bm{\mathbf{I}})/2$ with multiplicity accounted, and vise versa. As a result, $\eta_{s}(\bm{P}_{2})=\eta(\bm{P})/2$.
@@
In order to obtain the absolute spectral gap of $\bm{P}_{2}$, we need to study the singular value of the matrix $\bm{D}_{\mu_{2}}^{1/2}\bm{P}_{2}\bm{D}_{\mu_{2}}^{-1/2}$. Let $\bm{S}_{1}=\bm{D}_{\mu_{2}}^{1/2}{\bm{S}}\bm{D}_{\mu}^{-1/2}$ and $\bm{T}_{1}=\bm{D}_{\mu}^{1/2}{\bm{T}}\bm{D}_{\mu_{2}}^{-1/2}$. From Theorem 3.3 we know
@@
$\left({\bm{S}_{1}}^{\mathrm{T}}{\bm{S}_{1}}\right)(u,v)$ $=\mu(u)^{-1/2}\mu(v)^{-1/2}\sum\limits_{\bm{u}=(u_{1},u_{2})\in{\Omega}^{2}}{\bm{S}}(\bm{u},u){\mu}_{2}(\bm{u}){\bm{S}}(\bm{u},v)$ (3.30)
@@
$\left({\bm{T}_{1}}{\bm{T}_{1}}^{\mathrm{T}}\right)(u,v)$ $=\mu(u)^{1/2}\mu(v)^{1/2}\sum\limits_{\bm{u}=(u_{1},u_{2})\in{\Omega}^{2}}{\bm{T}}(u,\bm{u}){\mu}_{2}(\bm{u})^{-1}{\bm{T}}(v,\bm{u})$ (3.31)
@@
We thus deduce $\bm{S}_{1}^{\mathrm{T}}\bm{S}_{1}=\bm{T}_{1}\bm{T}_{1}^{\mathrm{T}}=\bm{\mathbf{I}}$. Since $\bm{S}_{1}\bm{T}_{1}\bm{T}_{1}^{\mathrm{T}}\bm{S}_{1}^{\mathrm{T}}$ and $\bm{S}_{1}^{\mathrm{T}}\bm{S}_{1}\bm{T}_{1}\bm{T}_{1}^{\mathrm{T}}=\bm{\mathbf{I}}$ share the same nonzero eigenvalues, we conclude that $1$ is the only nonzero singular value of $\bm{D}_{\mu_{2}}^{1/2}\bm{P}_{2}\bm{D}_{\mu_{2}}^{-1/2}$ and the multiplicity of $1$ is $d$. Thus $\lambda_{a}(\bm{P}_{2})=1$ and $\eta_{a}(\bm{P}_{2})=1-\lambda_{a}(\bm{P}_{2})=0$.
@@
Finally, when $\bm{P}$ is reversible, we have
@@
$\bm{K}^{2}$ $=\frac{1}{4}\left(\bm{S}_{1}\bm{T}_{1}\bm{S}_{1}\bm{T}_{1}+\bm{S}_{1}\bm{T}_{1}\bm{T}_{1}^{\mathrm{T}}\bm{S}_{1}^{\mathrm{T}}+\bm{T}_{1}^{\mathrm{T}}\bm{S}_{1}^{\mathrm{T}}\bm{S}_{1}\bm{T}_{1}+\bm{T}_{1}^{\mathrm{T}}\bm{S}_{1}^{\mathrm{T}}\bm{T}_{1}^{\mathrm{T}}\bm{S}_{1}^{\mathrm{T}}\right)$ (3.41)
@@
which implies $\eta_{s}(\bm{P}_{2})=\eta(\bm{P})/2$. This completes the proof. ∎
@@
$\tilde{\bm{T}}\tilde{\bm{S}}(u,u)$ $=\sum\limits_{\tilde{\bm{u}}\in\tilde{\Omega}^{2}}\tilde{\bm{T}}(u,\tilde{\bm{u}})\tilde{\bm{S}}(\tilde{\bm{u}},u)=\sum\limits_{\tilde{\bm{u}}\sim(u,v),v\in\Omega}\tilde{\bm{T}}(u,\tilde{\bm{u}})\tilde{\bm{S}}(\tilde{\bm{u}},u)$ (3.53)
@@ 4.1. Tail bounds of MLE and SCE
We will first derive tail bounds for the error of MLE and SCE using concentration inequalities for Markov chains. It is remarkable that most of the existing results about concentration inequalities for Markov chains typically require a nonzero absolute spectral gap of the transition matrix, which are inapplicable to $\bm{P}_{2}$ since $\eta_{a}(\bm{P}_{2})=0$. Therefore, we choose to adopt the following Bernstein-type inequality obtained in [8], which only requires a nonzero symmetric spectral gap of the transition matrix.
@@
Let $\{u_{k}\}_{k\geq 1}$ be a Markov chain on $\Omega$ generated by a transition kernel $\bm{P}$ with an invariant distribution $\mu$ and a symmetric spectral gap $\eta_{s}(\bm{P})>0$, and let $g:\Omega\rightarrow\mathbb{R}$ be a scalar function that satisfies $\mathbb{E}_{\mu}[g]=0$, $\mathbb{E}_{\mu}[g^{2}]\leq\sigma^{2}$, and $|g|\leq M$. It holds that for any $t>0$,
$\mathbb{P}\left\{\left|\frac{1}{n}\sum\limits_{j=1}^{n}g(u_{j})\right|\geq t\right\}\leq 2\|\nu/\mu\|_{\infty}\exp\left(\frac{-n\eta_{s}(\bm{P})t^{2}}{4M\sqrt{\left(2+5\eta_{s}(\bm{P})/\sqrt{2}\right)^{2}\sigma^{2}+t^{2}}}\right),$ (4.1)
@@
We then use Theorem 4.1 to derive non-asymptotic tail bounds for MLE.
@@
Next, for any $u,v\in\Omega$, we take $g=\mathbb{1}_{(u,v)}-\mu(u)p(u,v)$. Then it is easy to verify that $\mathbb{E}_{\mu}[g]=0$, $|g|\leq 1$, and $\mathbb{E}_{\mu}[g^{2}]\leq 1$. We therefore substitute $M=\sigma^{2}=1$, $\eta_{s}(\bm{P}_{2})\geq\eta_{s}(\bm{P})/2$ into Theorem 4.1 and deduce
$\mathbb{P}\left\{\left|r_{n}(u,v)-\mu(u)p(u,v)\right|\geq t\right\}\leq 2\|\nu_{2}/\mu_{2}\|_{\infty}\exp\left(\frac{-n\eta_{s}(\bm{P})t^{2}}{8\sqrt{\left(2+5\eta_{s}(\bm{P})/2\sqrt{2}\right)^{2}+t^{2}}}\right),$ (4.4)
where $\nu_{2}$ is the distribution of $\bm{u}_{1}$. Finally, for any $\bm{w}=(w_{1},w_{2})\in\Omega^{2}$, we have
@@
which implies $\|\nu_{2}/\mu_{2}\|_{\infty}=\|\nu/\mu\|_{\infty}$. The proof is completed. ∎
As for SCE, we choose to derive non-asymptotic tail bounds directly in the matrix form using matrix concentration inequalities. Recently, [16] established Hoeffding and Bernstein-type matrix concentration inequalities (in $l_{2}$ operator norm) for sums of Markov dependent random matrices, under the assumption that the transition matrix admits a nonzero absolute spectral gap. We present a simplified version of [16, Theorem 2.2] as follows.
@@
Let $\{u_{k}\}_{k\geq 1}$ be a Markov chain on $\Omega$ generated by a transition kernel $\bm{P}$ with an invariant distribution $\mu$ and an absolute spectral gap $\eta_{a}(\bm{P})>0$, and let $\bm{G}:\Omega\rightarrow\mathbb{H}_{d}$ be a matrix-valued function that satisfies $\mathbb{E}_{\mu}[\bm{G}]=0$, $\|\mathbb{E}_{\mu}[\bm{G}^{2}]\|\leq\sigma^{2}$, and $\|\bm{G}\|\leq M$. It holds that for any $t>0$,
$\mathbb{P}\left\{\left\|\frac{1}{n}\sum\limits_{j=1}^{n}\bm{G}(u_{j})\right\|\geq t\right\}\leq 2\|\nu/\mu\|_{\infty}d^{2-\frac{\pi}{4}}\exp\left(\frac{-n\eta_{a}(\bm{P})t^{2}}{32\pi^{-2}(2-\eta_{a}(\bm{P}))\sigma^{2}+256\pi^{-3}Mt}\right),$ (4.6)
@@
in Lemma 3.2, and then the almost surely convergence of $\bm{H}_{n}$ follows immediately. Next, we apply Theorem 4.2 with $\bm{G}=\tilde{\bm{F}}-\bm{D}_{\mu}\bm{P}$. We find that
$\left\|\bm{G}\right\|\leq\left\|\tilde{\bm{F}}\right\|+\left\|\bm{D}_{\mu}\bm{P}\right\|\leq 2,$ (4.10)
@@
$\left\|\mathbb{E}_{\tilde{\mu}_{2}}\left[\bm{G}^{2}\right]\right\|=\left\|\mathbb{E}_{\tilde{\mu}_{2}}\left[\tilde{\bm{F}}^{2}\right]-\left(\bm{D}_{\mu}\bm{P}\right)^{2}\right\|\leq\left\|\mathbb{E}_{\tilde{\mu}_{2}}\left[\tilde{\bm{F}}^{2}\right]\right\|+\left\|\bm{D}_{\mu}\bm{P}\right\|^{2}\leq 2.$ (4.11)
Substituting $M=\sigma^{2}=2$, $\eta_{a}(\bm{\tilde{P}_{2}})=\eta(\bm{P})/2$ into Theorem 4.2, we have
$\mathbb{P}\left\{\left\|\bm{H}_{n}-\bm{D}_{\mu}\bm{P}\right\|\geq t\right\}\leq 2\|\tilde{\nu}_{2}/\tilde{\mu}_{2}\|_{\infty}d^{2-\frac{\pi}{4}}\exp\left(\frac{-n\eta(\bm{P})t^{2}}{128\pi^{-2}(2-\eta(\bm{P})/2)+1024\pi^{-3}t}\right),$ (4.12)
@@
Therefore, we have $\|\tilde{\nu}_{2}/\tilde{\mu}_{2}\|_{\infty}=\|\nu/\mu\|_{\infty}$, which implies the claimed tail bound. ∎
Remark 4.3.
The reason we can establish matrix-form tail bounds for SCE is that $\tilde{\bm{P}}_{2}$ always admits a nonzero absolute spectral gap as long as $\bm{P}$ is irreducible. In contrast, the absolute spectral gap of $\bm{P}_{2}$ is always zero, which makes the result proposed in [16] inapplicable to MLE.

### v2

@@ 3.1.3. Invariant distributions
$\sum\limits_{\bm{v}\in\Omega^{2}}\mu_{2}(\bm{v})p_{2}(\bm{v},\bm{u})$ $=\sum\limits_{v\in\Omega}\mu_{2}\left(v,u_{1}\right)p\left(u_{1},u_{2}\right)$
@@ 3.3. Spectral properties of $\bm{P}_{2}$ and $\tilde{\bm{P}}_{2}$
In order to derive non-asymptotic error bounds for MLE and SCE, we need to study the spectral gaps of $\bm{P}_{2}$ and $\tilde{\bm{P}}_{2}$, respectively. To start, the next theorem focuses on the spectral properties of $\bm{P}_{2}$. We will show that $\bm{P}$ and $\bm{P}_{2}$ share the same nonzero eigenvalues with algebraic multiplicity accounted by decomposing $\bm{P}_{2}$ as the product of two low-rank matrices according to the construction of $\bm{P}_{2}$. We further show that the pseudo spectral gap, the symmetric spectral gap and the IP gap of $\bm{P}_{2}$ can all be bounded from below in terms of the corresponding spectral gaps of $\bm{P}$, however, the absolute spectral gap of $\bm{P}_{2}$ is always zero.
@@
    For any $k\in\mathbb{N}_{+}$, $\bm{P}^{k-1}$ and $\bm{P}_{2}^{k}$ share the same nonzero singular values with multiplicity accounted. As a result, $\eta_{a}(\bm{P}_{2})\equiv 0$, $\eta_{ps}(\bm{P}_{2})\geq\eta_{ps}(\bm{P})/2$.
@@
    $\eta_{s}(\bm{P}_{2})\geq\eta_{s}(\bm{P})/2>0$, and hence, $\mu_{2}$ is the unique invariant distribution of $\bm{P}_{2}$. Moreover, when $\bm{P}$ is reversible, any nonzero eigenvalue of $(\bm{P}_{2}+\bm{P}^{*}_{2})/2$ must be a nonzero eigenvalue of $(\bm{P}+\bm{\mathbf{I}})/2$ or $(\bm{P}-\bm{\mathbf{I}})/2$ with multiplicity accounted, and vise versa. As a result, $\eta_{s}(\bm{P}_{2})=\eta(\bm{P})/2$.
@@
    $\eta_{p}(\bm{P}_{2})\geq\eta_{p}(\bm{P})/(1+\eta_{p}(\bm{P}))>0$.
@@
In order to obtain the absolute spectral gap and pseudo spectral gap of $\bm{P}_{2}$, we need to study the singular value of the matrix ($\bm{D}_{\mu_{2}}^{1/2}\bm{P}_{2}\bm{D}_{\mu_{2}}^{-1/2})^{k}$. Let $\bm{S}_{1}=\bm{D}_{\mu_{2}}^{1/2}{\bm{S}}\bm{D}_{\mu}^{-1/2}$ and $\bm{T}_{1}=\bm{D}_{\mu}^{1/2}{\bm{T}}\bm{D}_{\mu_{2}}^{-1/2}$. From Theorem 3.3 we know
@@
$\left({\bm{S}_{1}}^{\mathrm{T}}{\bm{S}_{1}}\right)(u,v)$ $=\mu(u)^{-1/2}\mu(v)^{-1/2}\sum\limits_{\bm{u}=(u_{1},u_{2})\in{\Omega}^{2}}{\bm{S}}(\bm{u},u){\mu}_{2}(\bm{u}){\bm{S}}(\bm{u},v)$
@@
$\left({\bm{T}_{1}}{\bm{T}_{1}}^{\mathrm{T}}\right)(u,v)$ $=\mu(u)^{1/2}\mu(v)^{1/2}\sum\limits_{\bm{u}=(u_{1},u_{2})\in{\Omega}^{2}}{\bm{T}}(u,\bm{u}){\mu}_{2}(\bm{u})^{-1}{\bm{T}}(v,\bm{u})$
@@
We thus deduce $\bm{S}_{1}^{\mathrm{T}}\bm{S}_{1}=\bm{T}_{1}\bm{T}_{1}^{\mathrm{T}}=\bm{\mathbf{I}}$. Since $(\bm{S}_{1}\bm{T}_{1})^{k}(\bm{T}_{1}^{\mathrm{T}}\bm{S}_{1}^{\mathrm{T}})^{k}$ and $\bm{S}_{1}^{\mathrm{T}}\bm{S}_{1}(\bm{T}_{1}\bm{S}_{1})^{k-1}\bm{T}_{1}\bm{T}_{1}^{\mathrm{T}}(\bm{S}_{1}^{\mathrm{T}}\bm{T}_{1}^{\mathrm{T}})^{k-1}=\bm{D}_{\mu}^{1/2}\bm{P}^{k-1}\bm{D}_{\mu}^{-1/2}$ share the same nonzero eigenvalues for any $k\in\mathbb{N}_{+}$, we conclude that $\bm{P}^{k-1}$ and $\bm{P}_{2}^{k}$ share the same nonzero singular values. We therefore have
$\lambda_{a}\left(\bm{P}_{2}^{k}\right)=\lambda_{a}\left(\bm{P}^{k-1}\right).$
In particular, by taking $k=1$, we deduce $\lambda_{a}(\bm{P}_{2})=1$ and $\eta_{a}(\bm{P}_{2})=1-\lambda_{a}(\bm{P}_{2})=0$. By the definition of the pseudo spectral gap, we know that
$\eta_{ps}(\bm{P})=\sup_{k\in\mathbb{N}_{+}}\frac{1}{k}\left(1-\lambda_{a}^{2}\left(\bm{P}^{k}\right)\right).$
Suppose the above supremum is attained at $k_{0}\in[1,+\infty]$, we have
$\eta_{ps}(\bm{P}_{2})\geq\frac{1}{k_{0}+1}\left(1-\lambda_{a}^{2}\left(\bm{P}_{2}^{k_{0}+1}\right)\right)=\frac{1}{k_{0}+1}\left(1-\lambda_{a}^{2}\left(\bm{P}^{k_{0}}\right)\right)=\frac{k_{0}}{k_{0}+1}\eta_{ps}(\bm{P})\geq\frac{\eta_{ps}(\bm{P})}{2},$
which is the desired result.
@@
When $\bm{P}$ is reversible, we have
@@
$\bm{K}^{2}$ $=\frac{1}{4}\left(\bm{S}_{1}\bm{T}_{1}\bm{S}_{1}\bm{T}_{1}+\bm{S}_{1}\bm{T}_{1}\bm{T}_{1}^{\mathrm{T}}\bm{S}_{1}^{\mathrm{T}}+\bm{T}_{1}^{\mathrm{T}}\bm{S}_{1}^{\mathrm{T}}\bm{S}_{1}\bm{T}_{1}+\bm{T}_{1}^{\mathrm{T}}\bm{S}_{1}^{\mathrm{T}}\bm{T}_{1}^{\mathrm{T}}\bm{S}_{1}^{\mathrm{T}}\right)$
@@
which implies $\eta_{s}(\bm{P}_{2})=\eta(\bm{P})/2$.
Finally, we study the IP gap of $\bm{P}_{2}$. For any function $h\in L_{2,\mu_{2}}^{0}$, let $f:=(\bm{\mathbf{I}}-\bm{P}_{2})h$. Our goal is to bound $\|f\|_{\mu_{2}}/\|h\|_{\mu_{2}}$. Recall that $\{u_{k}\}_{k\geq 1}$ is the Markov chain generated by $\bm{P}$ and $\{\bm{u}_{k}\}_{k\geq 1}=\{(u_{k},u_{k+1})\}_{k\geq 1}$ is the Markov chain generated by $\bm{P}_{2}$. By the transition law of $\bm{P}_{2}$, for any $\bm{u}=(u,v)\in\Omega^{2}$,
$\bm{P}_{2}h(\bm{u})=\operatorname{\mathbb{E}}[h(\bm{u}_{2})|\bm{u}_{1}=\bm{u}]=\operatorname{\mathbb{E}}[h(v,u_{3})|u_{2}=v]=:h_{1}(v).$
we deduce that
$f(u,v)=h(u,v)-h_{1}(v),\qquad\text{for any }u,v\in\Omega.$
We then fix the first variable $u$ and take expectation with respect to the second variable $v$ according to the transition law $p(u,v)$ to get
$\operatorname{\mathbb{E}}[f(u,u_{2})|u_{1}=u]=\operatorname{\mathbb{E}}[h(u,u_{2})|u_{1}=u]-\operatorname{\mathbb{E}}[h_{1}(u_{2})|u_{1}=u]=(\bm{\mathbf{I}}-\bm{P})h_{1}(u).$
Since $\bm{P}$ admits an IP gap $\eta_{p}(\bm{P})>0$, we have
$\|h_{1}\|_{\mu}$ $\leq\frac{1}{\eta_{p}(\bm{P})}\|\operatorname{\mathbb{E}}[f(\cdot,u_{2})|u_{1}=\cdot]\|_{\mu}$
$=\frac{1}{\eta_{p}(\bm{P})}\left(\operatorname{\mathbb{E}}\left[\operatorname{\mathbb{E}}[f(u,u_{2})|u_{1}=u]^{2}|u\sim\mu\right]\right)^{1/2}$
$\leq\frac{1}{\eta_{p}(\bm{P})}\left(\operatorname{\mathbb{E}}\left[\operatorname{\mathbb{E}}[f^{2}(u,u_{2})|u_{1}=u]|u\sim\mu\right]\right)^{1/2}$
$=\frac{\|f\|_{\mu_{2}}}{\eta_{p}(\bm{P})}.$
As a result,
$\|f\|_{\mu_{2}}=\|h-h_{1}\|_{\mu_{2}}\geq\|h\|_{\mu_{2}}-\|h_{1}\|_{\mu_{2}}=\|h\|_{\mu_{2}}-\|h_{1}\|_{\mu}\geq\|h\|_{\mu_{2}}-\frac{\|f\|_{\mu_{2}}}{\eta_{p}(\bm{P})},$
which implies
$\frac{\|(\bm{\mathbf{I}}-\bm{P}_{2})h\|_{\mu_{2}}}{\|h\|_{\mu_{2}}}=\frac{\|f\|_{\mu_{2}}}{\|h\|_{\mu_{2}}}\geq\frac{\eta_{p}(\bm{P})}{1+\eta_{p}(\bm{P})}.$
This completes the proof.
@@
$\tilde{\bm{T}}\tilde{\bm{S}}(u,u)$ $=\sum\limits_{\tilde{\bm{u}}\in\tilde{\Omega}^{2}}\tilde{\bm{T}}(u,\tilde{\bm{u}})\tilde{\bm{S}}(\tilde{\bm{u}},u)=\sum\limits_{\tilde{\bm{u}}\sim(u,v),v\in\Omega}\tilde{\bm{T}}(u,\tilde{\bm{u}})\tilde{\bm{S}}(\tilde{\bm{u}},u)$
@@ 4.1. Tail bounds of MLE and SCE
We will first derive tail bounds for the error of MLE and SCE using concentration inequalities for Markov chains. Let us state the complete version of [10, Theorem 2.4], which is a Bernstein-type inequality for Markov sums that only requires a nonzero IP gap of the transition matrix.
@@
Let $\{u_{k}\}_{k\geq 1}$ be a Markov chain on $\Omega$ generated by a transition kernel $\bm{A}$ with an invariant distribution $\mu$ and an IP gap $\eta_{p}(\bm{A})>0$, and let $g:\Omega\rightarrow\mathbb{R}$ be a scalar function that satisfies $\mathbb{E}_{\mu}[g]=0$, $\mathbb{E}_{\mu}[g^{2}]\leq\sigma^{2}$, and $|g|\leq M$. It holds that for any $t>0$,
$\mathbb{P}\left\{\left|\frac{1}{n}\sum\limits_{j=1}^{n}g(u_{j})\right|\geq t\right\}\leq 2\|\nu/\mu\|_{\infty}\exp\left(\frac{-n\,\eta_{p}(\bm{A})\,t^{2}}{4M\sqrt{\left(2+6\eta_{p}(\bm{A})\right)^{2}\sigma^{2}+t^{2}}}\right),$ (4.1)
@@
In the above theorem, we use the symbol $\bm{A}$ for an arbitrary Markov transition kernel, so that it will not be confused with the particular transition matrix $\bm{P}$ in our context. Since $\eta_{p}(\bm{A})\geq\eta_{s}(\bm{A})\geq\eta_{a}(\bm{A})$ and $\eta_{p}(\bm{A})\geq\eta_{ps}(\bm{A})/2$, this concentration inequality also holds if the transition admits a nonzero absolute spectral gap, a nonzero pseudo spectral gap,or a nonzero symmetric spectral gap. We then use Theorem 4.1 to derive non-asymptotic tail bounds for MLE.
@@
Next, for any $u,v\in\Omega$, we take $g=\mathbb{1}_{(u,v)}-\mu(u)p(u,v)$. Then it is easy to verify that $\mathbb{E}_{\mu}[g]=0$, $|g|\leq 1$, and $\operatorname{Var}_{\mu}[g]\leq\mu_{2}(u,v)=\mu(u)p(u,v)$. Recall that $\bm{P}_{2}$ is the transition matrix of $\{\bm{u}_{k}\}_{k\geq 1}$. We can therefore apply Theorem 4.1 with $\bm{A}=\bm{P}_{2}$, $M=1$, $\sigma^{2}\leq\mu_{2}(u,v)$ to deduce
$\mathbb{P}\left\{\left|r_{n}(u,v)-\mu(u)p(u,v)\right|\geq t\right\}\leq 2\|\nu_{2}/\mu_{2}\|_{\infty}\exp\left(\frac{-n\,\eta_{p}(\bm{P}_{2})\,t^{2}}{4\sqrt{\left(2+6\eta_{p}(\bm{P}_{2})\right)^{2}\mu_{2}(u,v)+t^{2}}}\right),$ (4.4)
where $\nu_{2}$ is the distribution of $\bm{u}_{1}$. Note that for any $\bm{w}=(w_{1},w_{2})\in\Omega^{2}$, we have
@@
which implies $\|\nu_{2}/\mu_{2}\|_{\infty}=\|\nu/\mu\|_{\infty}$. Finally, we use the fact $\eta_{p}(\bm{P}_{2})\geq\eta_{p}(\bm{P})/(1+\eta_{p}(\bm{P}))$ to conclude the desired result. ∎
As for SCE, we choose to derive non-asymptotic tail bounds directly in the matrix form using matrix concentration inequalities. Recently, [19] established Hoeffding and Bernstein-type matrix concentration inequalities (in $l_{2}$ operator norm) for sums of Markov dependent random matrices, under the assumption that the transition matrix admits a nonzero absolute spectral gap. We present a simplified version of [19, Theorem 2.2] as follows.
@@
Let $\{u_{k}\}_{k\geq 1}$ be a Markov chain on $\Omega$ generated by a transition kernel $\bm{A}$ with an invariant distribution $\mu$ and an absolute spectral gap $\eta_{a}(\bm{A})>0$, and let $\bm{G}:\Omega\rightarrow\mathbb{H}_{d}$ be a matrix-valued function that satisfies $\mathbb{E}_{\mu}[\bm{G}]=0$, $\|\mathbb{E}_{\mu}[\bm{G}^{2}]\|\leq\sigma^{2}$, and $\|\bm{G}\|\leq M$. It holds that for any $t>0$,
$\mathbb{P}\left\{\left\|\frac{1}{n}\sum\limits_{j=1}^{n}\bm{G}(u_{j})\right\|\geq t\right\}\leq 2\|\nu/\mu\|_{\infty}d^{2-\frac{\pi}{4}}\exp\left(\frac{-n\,\eta_{a}(\bm{A})\,t^{2}}{32\pi^{-2}(2-\eta_{a}(\bm{A}))\sigma^{2}+256\pi^{-3}Mt}\right),$ (4.6)
@@
in Lemma 3.2, and then the almost surely convergence of $\bm{H}_{n}$ follows immediately.
Next, we find that
$\left\|\tilde{\bm{F}}-\bm{D}_{\mu}\bm{P}\right\|\leq\left\|\tilde{\bm{F}}\right\|+\left\|\bm{D}_{\mu}\bm{P}\right\|\leq 2,$ (4.10)
@@
$\left\|\mathbb{E}_{\tilde{\mu}_{2}}\left[\left(\tilde{\bm{F}}-\bm{D}_{\mu}\bm{P}\right)^{2}\right]\right\|=\left\|\mathbb{E}_{\tilde{\mu}_{2}}\left[\tilde{\bm{F}}^{2}\right]-\left(\bm{D}_{\mu}\bm{P}\right)^{2}\right\|\leq\left\|\mathbb{E}_{\tilde{\mu}_{2}}\left[\tilde{\bm{F}}^{2}\right]\right\|+\left\|\bm{D}_{\mu}\bm{P}\right\|^{2}\leq 2.$ (4.11)
We then apply Theorem 4.2 with $\bm{A}=\tilde{\bm{P}}_{2}$, $\bm{G}=\tilde{\bm{F}}-\bm{D}_{\mu}\bm{P}$, $M=\sigma^{2}=2$ to deduce that
$\mathbb{P}\left\{\left\|\bm{H}_{n}-\bm{D}_{\mu}\bm{P}\right\|\geq t\right\}\leq 2\|\tilde{\nu}_{2}/\tilde{\mu}_{2}\|_{\infty}d^{2-\frac{\pi}{4}}\exp\left(\frac{-n\,\eta_{a}(\tilde{\bm{P}}_{2})\,t^{2}}{64\pi^{-2}(2-\eta_{a}(\tilde{\bm{P}}_{2}))+512\pi^{-3}t}\right),$ (4.12)
@@
Therefore, we have $\|\tilde{\nu}_{2}/\tilde{\mu}_{2}\|_{\infty}=\|\nu/\mu\|_{\infty}$. Finally, we use the fact that $\eta_{a}(\bm{\tilde{P}_{2}})=\eta(\bm{P})/2$ to conclude the proof. ∎